OpenStudy (anonymous):

A set of equations is given below: Equation C: y = 4x + 8 Equation D: y = 4x + 2 Which of the following best describes the solution to the given set of equations? A. No solution B. One Solution C. Two solution D. Infinitely many solution

pooja195 (pooja195):

y = 4x + 8 <---e1 y = 4x + 2 <--e2 Both of these equations tell us what y equals we can use substitution to solve the problem Plug e1 into e2 (either way you will get the same answer) 4x+8=4x+2 Solve for x

OpenStudy (anonymous):

x=4?

pooja195 (pooja195):

How did you get that?

OpenStudy (anonymous):

I subtracted 4x and 4x and then I did 8/4

pooja195 (pooja195):

4x-4x=0

pooja195 (pooja195):

8=0x+2 subtract 2 from both sides 8-2=?

OpenStudy (prettygirl_shynice):

u got it

OpenStudy (anonymous):

so you don't divide ?

pooja195 (pooja195):

Well you really can't 6=0 How many solutions are there ?

OpenStudy (anonymous):

no solution?

pooja195 (pooja195):

Well done :)

OpenStudy (anonymous):

thank you so much

pooja195 (pooja195):

yw ^_^